(8/18) – (8/15) = 8/3
 (8/18) – (8/15) = 8/3
 (8/18) – (8/15) = 8/3

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The given statement isFalse.
Consider Left Hand Side (LHS), LCM of 18 and 15 = 90

Then,
(8/18) = (8 × 5)/(18 × 5) = 40/90
(8/15) = (8 × 6)/(15 × 6) = 48/90
Difference of two fractions (40/90)–(48/90)

= -8/90 Right Hand Side (RHS) = 8/3
By comparing LHS and RHS,

LHS≠RHS -8/90≠8/3



Thanks